A company has a net profit margin of 12.0%, asset turnover of 0.6, and a liabilities-to-asset ratio of 55.0 percent. What is the ROE?
Answer #1
D/A =0.55
E/A=1-D/A=1-0.55=0.45
Equity multiplier=1/(E/A)=1/0.45=2.2222
DUPONT
ROE = Net profit margin*total asset turnover*equity multiplier
ROE%=0.12*0.6*2.2222
ROE% = 16